The sum of the measures of two angles is 280 degrees. Three tomes the measure of one angle is 24 less than the measure of the other angle. What is the measure of each angle?

Answer by checkley77(12844)

X+Y=280 or X=280-Y
3X=Y-24
3(280-Y)=Y-24
840-3Y=Y-24
-3Y-Y=-24-840
-4Y=-864
Y=-864/-4
Y=216 DEGREES IS THE LARGER ANGLE.
280-216=64 DEGREES IS THE SMALLER ANGLE.
PROOF:
3*64=216-24
192=192